OpenStudy (anonymous):

Help! Please!!!!!!!!!! x-3 is a factor of f(x)=x^3+2+5 Which method is the most convenient? How can I determine the end behavior of that function? The process in finding the solutions for that function and how these solutions are related to the graph of that function?

OpenStudy (anonymous):

@campbell_st @Directrix

OpenStudy (campbell_st):

can you rewrite the equation

OpenStudy (anonymous):

f(x)=x^3+2+5 is my equation but it's asking the different methods determine if x-3 is a factor of my equation @campbell_st

OpenStudy (campbell_st):

so why is it written as \[f(x) = x^3 + 2 + 5 ~~~instead~ of~~~f(x) = x^3 + 7\]

OpenStudy (anonymous):

Because I need three or more terms

OpenStudy (campbell_st):

to determine if its a factor x - 3 as a factor then x = 3 is a zero... so find f(3) if it gives a zero, then you have x - 3 as a factor. its called the factor theorem... or remainder theorem

OpenStudy (anonymous):

Okay now How can I determine the end behavior of that function? The process in finding the solutions for that function and how these solutions are related to the graph of that function? @campbell_st

OpenStudy (campbell_st):

but based on the information, if every term is positive then (x -3) can't be a factor... so how are you going to find the 3 terms...?

OpenStudy (anonymous):

Can you make a new polynomial for me with a degree of 3 or higher and has 3 or more terms? @campbell_st

OpenStudy (campbell_st):

so is (x - 3) a factor...?

OpenStudy (anonymous):

yeah

OpenStudy (campbell_st):

ok... make it simple.... f(x) = x(x -3)^2 that is degree 3, with (x -3) as a factor... now distribute to get the equation in standard form.

OpenStudy (campbell_st):

so you were asked to create any degree 3 polynomial..with (x -3) as a factor and minimum of 3 terms

OpenStudy (anonymous):

yeah @campbell_st

OpenStudy (campbell_st):

ok... so distribute and you'll have a polynomial in standard form

OpenStudy (anonymous):

How do I do that? @campbell_st